OpenStudy (anonymous):

f(x) = x3 + 5x2 - 28x - 32 1 imaginary; 2 real 2 imaginary; 1 real 0 imaginary; 3 real 3 imaginary; 0 real

OpenStudy (anonymous):

I got d

ganeshie8 (ganeshie8):

hint : Any cubic polynomial will have atleast 1 real root

OpenStudy (anonymous):

Ive already found my answer is that right

ganeshie8 (ganeshie8):

hint2 : imaginary roots always come in pairs

ganeshie8 (ganeshie8):

did u take time to read my first hint

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohhhh

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

Having knowledge of those 2 hints, you can strike off few options. wat wud be they ?

OpenStudy (anonymous):

a and d

ganeshie8 (ganeshie8):

yes answer is either b or c

ganeshie8 (ganeshie8):

you familiar wid descarte's rule of signs ?

OpenStudy (anonymous):

um I dont think I know it by name, but if you told it to me I might idk.. hahaha

ganeshie8 (ganeshie8):

f(x) = x3 + 5x2 - 28x - 32

ganeshie8 (ganeshie8):

count the sign changes

OpenStudy (anonymous):

two

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

one

ganeshie8 (ganeshie8):

yes, so 1 positive real root exists.

ganeshie8 (ganeshie8):

Next, take f(-x)

OpenStudy (anonymous):

AYYYY thanks bro! its gotta be b

ganeshie8 (ganeshie8):

hey we're not finished yet, one more step is there to conlcude fully

OpenStudy (anonymous):

oh ok

ganeshie8 (ganeshie8):

you need to count sign changes in f(-x) also

OpenStudy (anonymous):

it 2

ganeshie8 (ganeshie8):

f(x) = x3 + 5x2 - 28x - 32 f(-x) = -x3 + 5x2 + 28x - 32

OpenStudy (anonymous):

thank you so much

ganeshie8 (ganeshie8):

it has 2 sign changes. so it can have 2 or 0 negative real roots

ganeshie8 (ganeshie8):

answer is C